# Universal Resources for QAOA and Quantum Annealing

### v1

@@ I Introduction
In this work we adopt a different strategy, which combines an empirical understanding of the states created by quantum optimization, with a study of emergent universal properties common to QAOA and QA. The first central result in our work is showing that the concentration of parameters in optimal QAOA circuits [23, 24, 17, 16, 25, 26] happens on universal QA passages of monotonically growing interaction and progressively decreasing mixer Hamiltonian. We show that these multilayer QAOA and QA protocols with said trajectories act as cooling protocols that create pseudo-Boltzmann probability distributions. This result goes beyond our previous work with single-layer circuits [27, 28], providing evidence that the non-adiabatic errors in QAOA and QA passages behave as thermal contributions, with the peculiarity that QAOA has an additional but vanishing error that originates in the Trotterized evolution. Our discovery enables the application of QAOA and QA protocols without previous optimization and their use as a cooling algorithm or simulators of partition functions with tuneable temperature. Finally, this study is supported by an interpretation of both multilayer QAOA and QA as paths in Hamiltonian space, deriving common metrics for the resources associated to these two protocols. We provide evidence that these resources—integrated angles, time, integrated interactions, and number of layers—scale very favorably with problem size and target state temperature.
The property of angle concentration in QAOA optimized trajectories is not a new discovery. This result has been demonstrated for MaxCut on regular graphs and the Sherringthon-Kirkpatrick model [23], and remarkable patterns have been repeatedly reported in multiple studies [24, 17, 16, 25, 26]. Furthermore, the transfer of angles between problems has also been used in the past among problems of similar families [29, 18, 30, 31], and there is evidence of smooth annealing paths constructed from QAOA optimal parameters [16]. However, our work is the first one to demonstrate the universality of these concentrated paths and their connection to continuous QA trajectories, while at the same time identifying the errors of those discretized annealing protocols as quasithermal distributions with tuneable properties, extending our physical interpretation of QAOA [27, 28] to both QAOA and QA simultaneously.
@@ II Quantum Approximate Optimization Algorithm
QAOA is a hybrid quantum-classical algorithm used for solving combinatorial optimization problems [3, 4]. The quantum part of this algorithm can be implemented on a programmable gate-based universal quantum computer. The QAOA algorithm involves preparing the initial state |$\ket{\psi_{+}}$, followed by applying the single-layer QAOA operator for $p$ rounds, given by:
|$\ket{\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}}=\prod_{i=1}^{p}\hat{\mathcal{Q}}_{i}(\gamma_{i},\theta_{i})\ket{\psi_{+}}.$ (1)
Initially, all qubits are prepared in a superposition state encompassing all possible solutions, as given by the application of $N$ Hadamard gates $\mathbb{H}$ onto the qubits’ ground states |$\ket{0}$
|$\ket{\psi_{+}}:=\ket{+}^{\otimes N}=\mathbb{H}^{\otimes N}\ket{0}^{\otimes N}$ (2)
For simplicity of notation and ease of understanding, we can gather all the $\gamma_{i}$ and $\theta_{i}$ values in $p$-dimensional vectors denoted by $\boldsymbol{\gamma}_{p}=(\gamma_{1},\ldots,\gamma_{p})$ and $\boldsymbol{\theta}_{p}=(\theta_{1},\ldots,\theta_{p})$. This state is evolved under the circuit defined by Eq. (1) with variational parameters $[\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}]$ that are tuned to minimize the expectation value of the energy. Thus, we have:
$\left\{\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}\right\}=\arg\min\left[\bra{\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}}\hat{H}_{\rm QSNet}\ket{\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}}\right].$ (3)
The set of variational parameters $\{\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}\}$ is estimated through a classical optimization process, and measuring the state |$\ket{\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}}$ on the computational basis provides an approximate solution to the optimization problem encoded by $\hat{H}_{{\rm QSNet}}$.
@@
where $\boldsymbol{J}$ and $\boldsymbol{h}$ are an $N$-by-$N$ square coupling matrix and a vector of $N$ coefficients respectively, with $N$ the number of qubits or spins, and $J_{nm}$, $h_{n}$ are the matrix and vector coefficients. The operators $\hat{\mathbb{X}}_{i}$ and ^$\hat{\mathbb{Z}}_{i}$ are matrices of order $2^{N}$ defined by the relations $\hat{\mathbb{X}}_{i}=\hat{\mathbb{I}}_{1}\otimes\ldots\otimes\hat{\mathbb{I}}_{i-1}\otimes\hat{\sigma}_{i}^{x}\otimes\hat{\mathbb{I}}_{i+1}\otimes\ldots\hat{\mathbb{I}}_{N}$ and ^$\hat{\mathbb{Z}}_{i}=\hat{\mathbb{I}}_{1}\otimes\ldots\otimes\hat{\mathbb{I}}_{i-1}\otimes\hat{\sigma}_{i}^{z}\otimes\hat{\mathbb{I}}_{i+1}\otimes\ldots\hat{\mathbb{I}}_{N}$. Here, $\hat{\sigma}_{i}^{\alpha}$ denotes the Pauli operator at site $i$-th along the direction $\alpha=x,z$, $\hat{\mathbb{I}}_{i}$ is the identity matrix of order 2 at site $i$. In Fig 1(a), we schematically represent the QAOA algorithm and a QSNet featuring a graph $\mathcal{G}_{m}$, where $m$ enumerates the number of numerical experimental implementations.
@@
where $\mathbf{{\boldsymbol{x}}}=(x_{1},\ldots,x_{N})$ with $x_{i}\in\{0,1\}$, and $\boldsymbol{Q}$ is an $N$-by-$N$ square symmetric matrix of real coefficients denoted by $Q_{nm}$. It is well known that these problems can be mapped to the Ising form (6). The binary variables are cast into ^$\hat{\mathbb{Z}}_{n}$, where ^$\hat{\mathbb{Z}}_{n}\ket{z_{n}}=z_{n}\ket{z_{n}}$, and $z_{n}=2x_{n}-1\in\{-1,+1\}$ denotes the spin value. The map produces a relationship between the variables of Eq.(7) with the QSNet Hamiltonian parameters given by $J_{n\neq m}=Q_{nm}$, $J_{nn}=0$, and $h_{n}=\sum_{m=1}^{N}Q_{nm}$. Our study is based on an extensive statistics of randomly generated QUBO problems defined on fully connected graphs, i.e. all off-diagonal elements of $\mathbf{Q}$ or $\mathbf{J}$ are nonzero. The nonzero values $Q_{nm}$ of each sample are randomly drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with mean $\mu=0$ and variance $\sigma^{2}=1$. We are interested in studying the main characteristics of individual experiments and the collective mean behavior across $m$ samples of QSNets.
@@ III Connection between QAOA and Quantum Annealing
$i\frac{d\Theta}{dt}\partial_{\Theta}\ket{\psi(\Theta)}=\left[\frac{d\Theta}{dt}\hat{H}_{x}+B(\Theta)\hat{H}_{{\rm QSNet}}\right]\ket{\psi(\Theta)}.$ (12)
@@
$i\partial_{\Theta}\ket{\psi(\Theta)}=\left[\hat{H}_{x}+\frac{d\Gamma}{d\Theta}\hat{H}_{{\rm QSNet}}\right]\ket{\psi(\Theta)}.$ (14)
@@ IV Multilayer QAOA studies
This section discusses the performance of optimized QAOA circuits over multiple random QUBO problems, sampled as per the description in Sec. II, with increasing number of qubits from $N=2$ to $20$, and up to $p=30$ layers of gates. This extensive study, has been performed over hundreds of problems—800 random problems for $N<20$ and 500 for $N=20$—, obtaining very good statistics for both the type of quantum states produced (c.f. Sec. IV.1) as well as for the angles that result from the optimization (c.f. Sec. IV.2).
@@ IV.2 Multilayer QAOA gives universal trajectories
In these curves, $\epsilon(p,N)$ is a small parameter characterizing the finite-size deviation of the passage from the radius 1 circle. As shown in Fig. 4, the QAOA angles distribute evenly along this limit trajectory, with very minor deviations. Indeed, for Fig. 4a, we find for $N=$ 8, 10, 12, 14, 16, 18, 20, decreasing corrections $\epsilon(N,30)=-0.030,-0.007,0.002,0.014,0.016,0.012,0.005$, with a standard deviation equal to $0.001$.
@@ IV.3 Integrated angles and computational cost
In Sec. III we introduced the integrated angles $\Theta_{\text{max}}$ and $\Gamma_{\text{max}}$ as plausible measures of the computational resources implied by both quantum annealing and the QAOA algorithm. Fig. 5 shows how these numbers increase with the number of layers for different problem sizes. As anticipated before, there is an exceptionally good collapse of the average values over the different problem sizes—illustrated here for $N=14,16,18,20$—with a smooth, monotonous increase of the resources with the number of layers. Remarkably, the integrated angles of the interaction term increase linearly with the number of layers $\Gamma_{\text{max}}\propto p$ and follows an algebraic decay with the system size, $\Gamma_{\text{max}}\propto N^{-1/2}$. The mixer contribution, on the other hand follows a trend that also becomes linear in the limit of large number of layers, $\Theta_{\text{max}}\propto p$, but is independent of the number of qubits, reinforcing the idea that $\Theta_{\text{max}}$ is a good measure of “time”. Note that in both cases the growth of the integrated angles with $p$ implies that they are a good proxy for the actual cost of implementing these circuits in a quantum computer, which also increases linearly with the number of layers.
@@ IV.6 Scaling of QAOA resources
The problem with the plot in Fig. 5 is that it only reflects the cost of making larger circuits, but not how much the solution “improves” with those resources. However, in order to make clear statements about the performance of the algorithm, we need to understand the growth of the cost with the solution quality and the problem size.
In quantum annealing, where the focus lies on preparing good ground states, a more canonical plot would be to study the time-to-solution, that is, what resources are needed to achieve a fixed overlap with the target state. Fortunately, in our study we already have a very general figure of merit—$\beta_{\text{high}}$, the temperature of the produced states (c.f. Fig. 3)—which is not a prefixed tolerance, but a property of the algorithm’s outcome.
@@
Figure 8: (Color online) Evolution of the effective temperature $\beta\equiv\beta_{\rm high}$ with QA resources $\Gamma_{\textnormal{max}}$ extrapolated from optimal QAOA circuits of $p=[1,...,30]$. We plot the results of 800 (500 for N=20) random QUBO instances of $N=$ 14, 16, 18, 20 spins (markers), together with the average of these results (solid lines), and the 90% percentile intervals (dashed lines).
@@ V Conclusions and outlook
In this work we have performed a physical study of the properties of optimized QAOA circuits and their connection to quantum annealing trajectories. We have established that there exists a formal connection between QAOA and QA, as both explore paths in Hamiltonian space. This connection becomes empirically exact, as the average angles of QAOA circuits collapse over universal trajectories that approach quantum annealing paths in the limit of infinite many layers. These optimal trajectories have been shown to produce, both in the QAOA and QA scenarios, similar pseudo-Boltzmann distributions—these are states whose probability distribution approximates an exponential $\exp(-\beta_{\text{high}}E)$ but have scrambled phases—. However, QAOA protocols exhibit an additional background noise that can be explained as the Trotter errors that result from approximating QA trajectories. In both cases, the coldest temperature of these distributions exhibits favorable scalings with respect to the problem size and the integrated resources, which in QAOA also correspond to the number of layers in the algorithm, $T\sim\mathcal{O}(p^{-1})$. However, both QAOA and QA paths can be used to produce hotter states simply by rescaling the Hamiltonian trajectories to shorter paths than the optimal one.
@@
Our study has also confirmed the utility of QAOA circuits and QA passages as simulators of partition functions and thermal distributions. Not only this thermal characterization of the global output distribution has shown a valid metric for the performance of QAOA algorithm [35], but we have shown that QAOA and QA protocols produce similar distributions. This seems to contradict the folklore picture that assumes QA to have a catastrophic performance for moderate speeds, suggesting instead that annealing errors behave like thermal noise, whose temperature can be controlled with the annealing speed. It would be both very interesting to verify this hypothesis with larger simulations and study the resilience of this picture in real hardware with other sources of noise. Indeed, while our performance metrics suggest a favorable behavior of QAOA on most instances of random problems that include spin glasses, it remains to confirm the nature of this scaling and its robustness in practical scenarios.
Finally, the utility of QAOA for thermal state simulation has applications beyond the approximation of NP-hard problems. The sampling of pseudo-Boltzmann distributions could be a useful subroutine for physical simulation, combinatorial optimization, or machine learning. In this sense, it would be interesting to verify the tunability of said simulators in real hardware, as well as their extension to other types of ansätze or Hamiltonians in line with recent results [37].
@@ Acknowledgements.
This work has been supported by European Commission FET Open project AVaQus Grant Agreement No. 899561, Comunidad de Madrid Sinergicos 2020 project NanoQuCo-CM (Y2020/TCS- 6545), CSIC Quantum Technologies Platform PTI-001, Spanish project PID2021-127968NBI00 funded by MICIU/AEI/10.13039/501100011033, and the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda.

### v2

@@ I Introduction
In this work we adopt a different strategy, which combines an empirical understanding of the states created by quantum optimization, with a study of emergent universal properties common to QAOA and QA. The first central result in our work is showing that the concentration of parameters in optimal QAOA circuits [23, 24, 17, 16, 25, 26] happens on universal QA passages of monotonically growing interaction and progressively decreasing mixer Hamiltonian. We show that these multilayer QAOA and QA protocols with said trajectories act as cooling protocols that create pseudo-Boltzmann probability distributions - i.e., distributions that fluctuate around the Boltzmann distribution in energy space. This result goes beyond our previous work with single-layer circuits [27, 28], providing evidence that the non-adiabatic errors in QAOA and QA passages behave as thermal contributions, with the peculiarity that QAOA has an additional but vanishing error that originates in the Trotterized evolution. Our discovery enables the application of QAOA and QA protocols without previous optimization and their use as a cooling algorithm or simulators of partition functions with tuneable temperature. Finally, this study is supported by an interpretation of both multilayer QAOA and QA as paths in Hamiltonian space, deriving common metrics for the resources associated to these two protocols. We provide evidence that these resources—integrated angles, time, integrated interactions, and number of layers—scale very favorably with problem size and target state temperature.
The property of angle concentration in QAOA optimized trajectories is not a new discovery. This result has been demonstrated for MaxCut on regular graphs and the Sherrington-Kirkpatrick model [23], and remarkable patterns have been repeatedly reported in multiple studies [24, 17, 16, 25, 26]. Furthermore, the transfer of angles between problems has also been used in the past among problems of similar families [29, 18, 30, 31], and there is evidence of smooth annealing paths constructed from QAOA optimal parameters [16]. However, our work is the first one to demonstrate the universality of these concentrated paths - universality defined over the ensembles of dense spin glasses - and their connection to continuous QA trajectories, while at the same time identifying the errors of those discretized annealing protocols as quasithermal distributions with tuneable properties, extending our physical interpretation of QAOA [27, 28] to both QAOA and QA simultaneously.
@@ II Quantum Approximate Optimization Algorithm
[…]
For simplicity of notation and ease of understanding, we can gather all the $\gamma_{i}$ and $\theta_{i}$ values in $p$-dimensional vectors denoted by $\boldsymbol{\gamma}_{p}=(\gamma_{1},\ldots,\gamma_{p})$ and $\boldsymbol{\theta}_{p}=(\theta_{1},\ldots,\theta_{p})$. This state is evolved under the circuit defined by Eq. (1) with variational parameters $\boldsymbol{\gamma}_{p},\boldsymbol{\theta}_{p}]$ that are tuned to minimize the expectation value of the energy. Thus, we have:
[…]
@@
where $\boldsymbol{J}$ and $\boldsymbol{h}$ are an NN-by-NN square coupling matrix and a vector of $N$ coefficients respectively, with $N$ the number of qubits or spins, and $J_{nm}$, $h_{n}$ are the matrix and vector coefficients. The operators $\hat{\mathbb{X}}_{i}$ and $\hat{\mathbb{Z}}_{i}$ are matrices of order $2^{N}$ defined by the relations $\hat{\mathbb{X}}_{i}=\hat{\mathbb{I}}_{1}\otimes\ldots\otimes\hat{\mathbb{I}}_{i-1}\otimes\hat{\sigma}_{i}^{x}\otimes\hat{\mathbb{I}}_{i+1}\otimes\ldots\hat{\mathbb{I}}_{N}$ and $\hat{\mathbb{Z}}_{i}=\hat{\mathbb{I}}_{1}\otimes\ldots\otimes\hat{\mathbb{I}}_{i-1}\otimes\hat{\sigma}_{i}^{z}\otimes\hat{\mathbb{I}}_{i+1}\otimes\ldots\hat{\mathbb{I}}_{N}$. Here, $\hat{\sigma}_{i}^{\alpha}$ denotes the Pauli operator at site $i$-th along the direction $\alpha=x,z$, $\hat{\mathbb{I}}_{i}$ is the identity matrix of order 2 at site $i$. In Fig 1(a), we schematically represent the QAOA algorithm and a QSNet featuring a graph $\mathcal{G}_{m}$, where $m$ enumerates the number of numerical experimental implementations.
@@
[…]
@@ III Connection between QAOA and Quantum Annealing
[…]
@@
[…]
@@ IV Multilayer QAOA studies
This section discusses the performance of optimized QAOA circuits over multiple random QUBO problems, sampled as per the description in Sec. II, with increasing number of qubits from $N=2$ to $20$, and up to $p=30$ layers of gates. This extensive study, has been performed over hundreds of problems—800 random problems for $N<20$ and 500 for N=20N=20—, obtaining very good statistics for both the type of quantum states produced (c.f. Sec. IV.1) as well as for the angles that result from the optimization (c.f. Sec. IV.2).
@@ IV.2 Multilayer QAOA gives universal trajectories
In these curves, $\epsilon(p,N)$ is a small parameter characterizing the finite-size deviation of the passage from the radius 1 circle. As shown in Fig. 4, the QAOA angles distribute evenly along this limit trajectory, with very minor deviations. Indeed, for Fig. 4a, we find for $N=$ 8, 10, 12, 14, 16, 18, 20, decreasing corrections $\epsilon(N,30)=-0.030,-0.007,0.002,0.014,0.016,0.012,0.005$, with a standard deviation equal to $0.001$. In Appendix B, we show how these results also extend to other models beyond QUBO, such as the maximum cut.
@@ IV.3 Integrated angles and computational cost
In Sec. III we introduced the integrated angles $\Theta_{\text{max}}$ and $\Gamma_{\text{max}}$ as plausible measures of the computational resources implied by both quantum annealing and the QAOA algorithm. Fig. 5 shows how these numbers increase with the number of layers for different problem sizes. As anticipated before, there is an exceptionally good collapse of the average values over the different problem sizes—illustrated here for N=14,16,18,20N=14,16,18,20—with a smooth, monotonous increase of the resources with the number of layers. Remarkably, the integrated angles of the interaction term increase linearly with the number of layers $\Gamma_{\text{max}}\propto p$ and follows an algebraic decay with the system size, $\Gamma_{\text{max}}\propto N^{-1/2}$. The mixer contribution, on the other hand follows a trend that also becomes linear in the limit of large number of layers, $\Theta_{\text{max}}\propto p$, but is independent of the number of qubits, reinforcing the idea that $\Theta_{\text{max}}$ is a good measure of “time”. Note that in both cases the growth of the integrated angles with $p$ implies that they are a good proxy for the actual cost of implementing these circuits in a quantum computer, which also increases linearly with the number of layers.
@@ IV.6 Scaling of QAOA resources
The problem with the plot in Fig. 5 is that it only reflects the cost of making larger circuits, but not how much the solution “improves” with those resources. However, in order to make clear statements about the performance of the algorithm, we need to understand the growth of the cost with the solution quality and the problem size. From a computational perspective, scaling refers to the relationship between the resource requirements of an algorithm and the growth of the problem size, as well as its impact on performance.
[…]
@@
Figure 8: (Color online) Evolution of the effective temperature $\beta\equiv\beta_{\rm high}$ with QA resources $\Gamma_{\textnormal{max}}$ extrapolated from optimal QAOA circuits of p=[1,,30]p=[1,...,30]. We plot the results of 800 (500 for N=20) random QUBO instances of $N=$ 14, 16, 18, 20 spins (markers), together with the average of these results (solid lines), and the 90% percentile intervals (dashed lines).
@@ V Conclusions and outlook
In this work we have performed a physical study of the properties of optimized QAOA circuits and their connection to quantum annealing trajectories. We have established that there exists a formal connection between QAOA and QA, as both explore paths in Hamiltonian space. This connection becomes empirically exact, as the average angles of QAOA circuits collapse over universal trajectories - defined on the family of dense QUBO problems - that approach quantum annealing paths in the limit of infinite many layers. These optimal trajectories have been shown to produce, both in the QAOA and QA scenarios, similar pseudo-Boltzmann distributions—these are states whose probability distribution approximates an exponential $\exp(-\beta_{\text{high}}E)$ but have scrambled phases—. However, QAOA protocols exhibit an additional background noise that can be explained as the Trotter errors that result from approximating QA trajectories. In both cases, the coldest temperature of these distributions exhibits favorable scalings with respect to the problem size and the integrated resources, which in QAOA also correspond to the number of layers in the algorithm, $T\sim\mathcal{O}(p^{-1})$. However, both QAOA and QA paths can be used to produce hotter states simply by rescaling the Hamiltonian trajectories to shorter paths than the optimal one.
@@
Our study has also confirmed the utility of QAOA circuits and QA passages as simulators of partition functions and thermal distributions. Not only this thermal characterization of the global output distribution has shown a valid metric for the performance of QAOA algorithm [35], but we have shown that QAOA and QA protocols produce similar distributions. This seems to contradict the folklore picture that assumes QA to have a catastrophic performance for moderate speeds, suggesting instead that annealing errors behave like thermal noise, whose temperature can be controlled with the annealing speed. It would be both very interesting to verify this hypothesis with larger simulations and study the resilience of this picture in real hardware with other sources of noise. Indeed, while our performance metrics suggest a favorable behavior of QAOA on most instances of random problems that include spin glasses, it remains to confirm the nature of this scaling and its robustness in practical scenarios, including a broader set of problems with structured and sparse connectivity.
Finally, the utility of QAOA for thermal state simulation has applications beyond the approximation of NP-hard problems. The sampling of pseudo-Boltzmann distributions could be a useful subroutine for physical simulation, combinatorial optimization, or machine learning. In this sense, it would be interesting to verify the tunability of said simulators in real hardware, the comparison to alternative methods, as well as their extension to other types of ansätze or Hamiltonians in line with recent results [37].
@@ Acknowledgements.
This work has been supported by European Commission FET Open project AVaQus Grant Agreement No. 899561, Comunidad de Madrid Sinergicos 2020 project NanoQuCo-CM (Y2020/TCS- 6545), CSIC Quantum Technologies Platform PTI-001, Spanish project PID2021-127968NBI00 funded by MICIU/AEI/10.13039/501100011033, and the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda. FJGR acknowledges financial support from the Spanish Government via the project PID2024-161371NB-C21 (MCIU/AEI/FEDER, EU).
Appendix A Numerical details
In this appendix we provide additional details about the numerical experiments that support the results reported in the manuscript.
All quantum circuits were simulated using exact state-vector calculations under zero-noise conditions. The optimization of the variational parameters involved in the numerical results of QAOA at any depth was performed using the L-BFGS-B method implemented in scipy [38]. The gradients were calculated analytically, and the convergence criteria established that the optimization stops when
$\frac{f^{k}-f^{k+1}}{\max(|f^{k}|,|f^{k+1}|,1)}\leq f_{tol}=2\cdot 10^{-9},$ (32)
or
$\max{|\textnormal{proj(}g_{i})|\;i=1,...,n}<=10^{-10},$ (33)
with $f^{k}$ the value of the cost function at iteration $k$ and $\textnormal{proj(}g_{i})$ the i-th component of the projected gradient. The parameters of a QAOA circuit with $p$ layers are initialized using the optimal parameters of p1p-1,
$\theta^{p}_{l}=\Theta^{p-1}_{\max}/p\;\;\;,\;\;\;\gamma^{p}_{l}=\Gamma^{p-1}_{\max}/p,$ (34)
for $l=\{1,...,p\}$, with $\theta^{1}=\pi/3$ and $\gamma^{1}=0.1$. This heuristic was tested in different scenarios and provided good results in practice.
As for the simulation of continuous quantum annealing passages (see Fig.7), we numerically integrated the system of ordinary differential equations given by the normalized continuous Schrödinger equation between $\Theta=0$ and $\Theta=1$,
[…]
The integration method used was the explicit Runge-Kutta method of order 5(4) implemented in scipy [38] with a maximum step size equal to $10^{-5}$.
Appendix B Universal trajectories on MaxCut
Figure 9: (Color online) Annealing trajectories defined by the rescaled optimal QAOA angles $\left(\overline{\Theta}_{n}/\overline{\Theta}_{\textnormal{max}},\overline{\Gamma}_{n}/\overline{\Gamma}_{\textnormal{max}}\right)$. (a) Average trajectories fitted to Eq. (27) for $N=$ 8, 10, 12, 14, 16, 18, and 20 qubits, computed from $p=30$ QAOA angles on 800 MaxCut instances (500 for $N=$ 20). We also plot the average angles for $N=20$ qubits as circles. (a) Average rescaled optimal angles for QAOA circuits with $N=18$ spins and increasing number of layers $p=$ 5, 10, 15, 20, 25, and 30 layers, computed over 800 MaxCut instances, together with a numerical fit to Eq. 27 for $p=30$ layers (solid line).
The figures presented in the main text show universal behaviors of randomly generated QUBO problems (7). In this appendix, we extend the scope of the results to a different family of NP-hard binary combinatorial optimization problems, the maximum cut problem (MaxCut). MaxCut is defined by the following classical energy:
$E_{{\rm MaxCut}}({\boldsymbol{x}})=-2\sum_{n,m=1}^{N}x_{n}Q_{n,m}(1-x_{m}).$ (36)
where $\mathbf{{\boldsymbol{x}}}=(x_{1},\ldots,x_{N})$ with $x_{i}\in\{0,1\}$, and $\boldsymbol{Q}$ is an NN-by-NN square symmetric matrix of real coefficients denoted by $Q_{nm}$. As in QUBO instances, we consider fully connected graphs, i.e. all off-diagonal elements of $\mathbf{Q}$ or $\mathbf{J}$ are nonzero values randomly drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with mean $\mu=0$ and variance $\sigma^{2}=1$. Although MaxCut (36) can be formulated as a QUBO (7) by allowing diagonal elements in the Q matrix, MaxCut exhibits a global symmetry $\mathbb{Z}_{2}$ generally absent in QUBO instances.
Fig.9 shows that the optimal QAOA angles on MaxCut instances collapse into universal curves. As was the case for the QUBO problem class, these curves define annealing paths described by deformed radius 1 circles given by equation (27). Specifically, we find small corrections $\epsilon(N,p=30)=-0.026\pm 0.001,-0.019\pm 0.001,-0.017\pm 0.001,-0.019\pm 0.001,-0.022\pm 0.002,-0.031\pm 0.002,-0.031\pm 0.002$ for $N=$ 8, 10, 12, 14, 16, 18, 20, respectively.